If the length of a rectangle is 2cm. longer than the width, and the rectangle's diagonal is the square root of 11 cm. Find the width of the rectangle.
Answer by stanbon(75887)

If the length of a rectangle is 2cm. longer than the width, and the rectangle's diagonal is the square root of 11 cm. Find the width of the rectangle.
----
Let the width be "w".
Then the length is "w+2".
---
Use Pythagoras
w^2 + (w+2)^2 = (sqrt(11))^2
----
w^2 + w^2 + 4w + 4 = 11
2w^2 + 4w - 7 = 0
----
w = [-4 +- sqrt(16 - 4*2*-7)]/4
----
w = [-4+- sqrt(72)]/4
----
w = [-4 +-6sqrt(2)]/4
----
Positive solution:
w = [-4 + 6sqrt(2)]
width = 1.12 cm
length = 3.12 cm
---------------------
Cheers,
Stan H.
================